Matthew is married and lives in Pennsylvania, which has a state income tax rate of 3.07%. He earns an annual salary of $47,530. Calculate his monthly take-home pay.
Answer by math_tutor2020(3816)


$47,530 is the gross annual salary before taxes are deducted.

3.07% = 3.07/100 = 0.0307
3.07% of $47,530 = 0.0307*47530 = 1,459.171 = 1,459.17

Matthew is charged $1,459.17 in income tax.
His annual take-home pay is 47,530 - 1,459.17 = 46,070.83
This does not account for federal income taxes or other taxes such as FICA.
However, since your teacher didn't mention those other taxes, then I'm assuming s/he only is considering state taxes.

Lastly, divide by 12 to get the monthly take-home pay.
(46,070.83)/12 = 3,839.23583333333
This rounds to $3,839.24

---------------------------------------------------------------
Here's another way to get the answer

The state wants 3.07% of his annual salary.
He keeps the remaining 100% - 3.07% = 96.93%

96.93% of 47,530 = 0.9693*47530 = 46,070.829 = 46,070.83
We get the same annual take-home pay as before, which will lead to the same monthly take-home pay.

---------------------------------------------------------------

Answer: $3,839.24